# Towards Efficient Training of Graph Neural Networks: A Multiscale Approach

### v2

@@ Abstract
Graph Neural Networks (GNNs) have emerged as a powerful tool for learning and inferring from graph-structured data, and are widely used in a variety of applications, often considering large amounts of data and large graphs. However, training on such data requires large memory and extensive computations. In this paper, we introduce a novel framework for efficient multiscale training of GNNs, designed to integrate information across multiscale representations of a graph. Our approach leverages a hierarchical graph representation, taking advantage of coarse graph scales in the training process, where each coarse scale graph has fewer nodes and edges. Based on this approach, we propose a suite of GNN training methods: such as coarse-to-fine, sub-to-full, and multiscale gradient computation. We demonstrate the effectiveness of our methods on various datasets and learning tasks.
@@ 1 Introduction
Large-scale graph networks arise in diverse fields such as social network analysis Kipf & Welling (2016); Defferrard et al. (2016), recommendation systems Tang et al. (2018), and bioinformatics Jumper et al. (2021), where relationships between entities are crucial to understanding system behavior Abadal et al. (2021). These problems involve the processing of large graph data sets necessitating scalable solutions. Some examples of those problems include identifying community structures in social networks Hamilton et al. (2017), optimizing traffic flow in smart cities Geng et al. (2019), or studying protein-protein interactions in biology Xu et al. (2024). In all these problems, efficient processing of graph data is required for practical reasons.
In recent years, many Graph Neural Networks (GNNs) algorithms have been proposed to solve large graph problems, such as GraphSage Hamilton et al. (2017). However, despite their success, when the involved graph is of a large scale, with many nodes and edges, it takes a long time and many resources to train the network Galmés et al. (2021).
@@
Figure 1: Coarse graph using random pooling. The Orange nodes in the fine graph are selected to the coarse graph.
(a) Fine original graph
(b) Coarse graph
Figure 2: Coarse graph using Topk pooling, meaning nodes with highest degrees are chosen. Orange nodes in the fine graph are the nodes selected to the coarse graph.
(a) Fine original graph
(b) Coarse graph
Figure 3: Coarse graph using subgraph pooling. Orange nodes in the fine graph are the nodes selected to the coarse graph, and the red node is the root node for the ego-network.
@@
We demonstrate the generation of a coarse graph using random pooling in Figure 1 and Topk pooling (which is based on the nodes with high degrees in the graph) in Figure 2. As for choosing a subgraph, this can be implemented using node-based selection methods, such as ego-networks (Gupta et al., 2014) or node marking and deletion (Frasca et al., 2022), which offer straightforward ways to extract subgraphs based on specific nodes of interest. We show an example of subgraph coarsening in Figure 3.
@@
While coarsening and subgraph methods have been in use for various purposes, from improved architectures Gao & Ji (2019), to enhanced expressiveness Bevilacqua et al. (2022; 2024), to the best of our knowledge, they were not considered in the context of efficient training of GNNs. Therefore, in this work, we explore three novel training mechanisms to design surrogate processes that approximate the original training problem with a fraction of the computational cost, using a multiscale approach. All the methods attempt to use the full graph with $N$ nodes as little as possible, and focus on training the network using coarse-scale graphs, thus reducing computational cost. The three algorithms, detailed in Section 4, are as follows:
  • Coarse-to-Fine. A coarse-to-fine algorithm approximates the optimization problem on smaller, pooled graphs, reducing costs. If the coarse-scale solution matches the fine-scale one, it can serve as an effective initialization for the fine-scale optimization. Since gradient-based methods converge faster when initialized near the solution (Arora et al., 2018; Boyd & Vandenberghe, 2004), this approach reduces the overall number of iterations required.
  • Sub-to-Full. This approach leverages a sequence of subgraphs that progressively grow to the original fine-scale graph. Like coarsening, each subgraph is significantly smaller, enabling faster optimization. Subgraphs can be generated based on node distances (if available) or ego-network methods Gupta et al. (2014).
@@
The main contribution of this paper is the design of novel graph multiscale optimization algorithms to obtain computationally efficient GNN training methods that retain the performance of standard gradient-based training of GNN. To demonstrate their effectiveness, we benchmark the obtained performance and reduce computational costs on multiple graph learning tasks, from transductive node classification to inductive tasks like shape segmentation.
@@ 2 Related Work
Graph pooling can be performed in various ways, as reviewed in Liu et al. (2022). Popular pooling techniques include DiffPool (Ying et al., 2018) MaxCutPool (Abate & Bianchi, 2024), graph wavelet compression Eliasof et al. (2023), as well as Independent-Set Pooling (Stanovic et al., 2024), all aiming to produce a coarsened a graph that bears similar properties to the original graph. Wang et al. (2020) and Wang et al. (2024) introduce the concept of subgraph pooling to GNNs. Bianchi & Lachi (2023) study the expressive power of pooling in GNNs. Here, we use graph pooling mechanisms, as illustrated in Figure 1Figure 3, as core components for efficient training.
@@ 3 Preliminaries and Notations
We now provide related graph-learning notations and preliminaries that we will use throughout this paper.
Let us consider a graph defined by ${\cal G=(\cal V,\cal E)}$ where $\cal V$ is a set of $n$ nodes and $\cal E$ is a set of $m$ edges. Alternatively, a graph can be represented by its adjacency matrix ${\bf A}$, where the $(i,j)$-th entry is 1 if an edge between nodes $i$ and $j$ exists, and 0 otherwise. We assume that each node is associated with a feature vector with $c$ channels as ${\bf x}_{i}\in\mathbb{R}^{c}$, and ${\bf X}=[{\bf x}_{0}\ldots,{\bf x}_{n-1}]\in\mathbb{R}^{n\times c}$ as the node features matrix. We denote by ${\bf y}_{i}\in\mathbb{R}^{d}$ the $i$-th node label, and ${\bf Y}=[{\bf y}_{0},\ldots,{\bf y}_{n-1}]\in\mathbb{R}^{n\times d}$ the label matrix.
@@
For instance, in the well-known Graph Convolution Network (GCN) (Kipf & Welling, 2016), each layer is defined using
Equation (1), such that ${\bf X}^{(l+1)}$ is the output of the $l$-th layer and serves as the input for the $(l+1)$-th layer. For a network with $L$ layers, we denote by ${\boldsymbol{\theta}}=\{{\bf W}^{(1)},\ldots,{\bf W}^{(L)}\}$ the collection of all learnable weight matrices within the network.
@@ 4 Efficient Training Framework for GNNs
${\bf Y}_{r}\approx f({\bf X}_{r},{{\cal G}_{r}};{\boldsymbol{\theta}}),\quad r=1,\ldots,R,$ (3)
where $f$ is a GNN and ${\cal G}_{r}$ is a graph with $n_{r}$ nodes, where $n=n_{1}>n_{2}>\ldots>n_{R}$. Each graph ${\cal G}_{r}$ is obtained by pooling the fine graph $\mathcal{G}$, using a smaller set of nodes compared to the previous (finer) graph $\mathcal{G}_{r-1}$.
${\bf X}_{r}$ is the node features matrix of the graph $\mathcal{G}_{r}$, and ${\bf Y}_{r}$ is the corresponding label matrix. In what follows, we explore different ways to obtain ${\cal G}_{r},{\bf X}_{r}$ and ${\bf Y}_{r}$ from the original resolution variables.
@@ 4.1 Coarse-to-Fine Training
While Algorithm 1 is relatively simple, it requires obtaining the $r$-th surrogate problem by coarsening the graph, as well as stopping a criterion for the $r$-th optimization problem. Generally, in our coarse-to-fine method, we use the following graph coarsening scheme:
@@
where ${\bf P}_{r}\in\mathbb{R}^{n\times n_{r}}$ is a binary injection matrix that maps the fine graph with $n$ nodes to a smaller graph with $n_{r}$ nodes. We use ${\bf X}_{r}={\bf P}_{r}^{T}{\bf X}$ as the coarse graph node features matrix and ${\bf Y}_{r}={\bf P}_{r}^{T}{\bf Y}$ as the label matrix of the coarse graph.
Choosing the non-zero entries of the matrix ${\bf P}_{r}$, which are the nodes that remain in the graph after pooling, can be implemented in various ways. For example, ${\bf P}_{r}$ can be obtained by sampling random nodes of the fine graph $\mathcal{G}$, as illustrated in Figure 1, or with some desired pooling logic, such as choosing nodes of the highest degree in the graph, which is a method often described as $Topk$ pooling (Huang et al., 2015) and we present an illustration of it in Figure 2. The power $p$ in Equation (5) controls the connectivity of the graph, where $p=1$ is the original graph adjacency matrix, and using a value $p>1$ will enhance the connectivity of the graph.
In Algorithm 1, we apply GNN layers such as in Equation (1), using the coarsened feature matrix ${\bf X}_{r}$ and its corresponding adjacency matrix ${\bf A}_{r}$, that is based on a desired pooling ratio $\rho$. We note that there is no re-initialization of the weight matrix ${\bf W}$ when transferring from coarse grids to finer ones, allowing a continuous learning process along the steps of the algorithm. As we show later in our experiments in Section 6, our coarse-to-fine training approach, summarized in Algorithm 1, is not restricted to a specific GNN architecture or a pooling technique.
@@ 4.2 Sub-to-Full Training
A subgraph can be defined in a few possible manners. One is by choosing a root node $j$, and generating a subgraph centered around that node. If the graph $\cal{G}$ has some geometric features, meaning each input node
corresponds to a point in the Euclidean space $\mathbb{R}^{d}$, one might consider constructing a subgraph with $n_{r}$ nodes to be the $n_{r}$ nodes that are closest to the chosen center node in terms of Euclidean distance norm. Otherwise, we can consider a subgraph created using an ego-network Gupta et al. (2014) of $k-hop$ steps, such that we consider all neighbors that are $k$ steps away from the central node. An illustration of this method is presented in Figure 3. The number of hops, $k$, is a hyperparameter of the algorithm process, and we increase it as we get from coarse levels to fine ones within the series of graph $\{\mathcal{G}_{r}\}_{r=1}^{R}$.
@@ 4.3 Multiscale Gradients Computation
${\mathbb{E}}\,\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})={\mathbb{E}}\,\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})+{\mathbb{E}}\,\left(\mathcal{L}^{(R-1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})\right)+\ldots+{\mathbb{E}}\,\left(\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(2)}_{s}({\boldsymbol{\theta}})\right),$ (8)
@@
${\mathbb{E}}\,\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})\approx\textstyle{\frac{1}{M_{R}}\sum\mathcal{L}^{(R)}_{j}({\boldsymbol{\theta}})}+\textstyle{\frac{1}{M_{R-1}}\sum\left(\mathcal{L}^{(R-1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})\right)}+\ldots+\textstyle{\frac{1}{M_{1}}\sum\left(\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(2)}_{s}({\boldsymbol{\theta}})\right)},$ (10)
where $M_{R}>M_{R-1}>M_{1}$. This implies that for the computation of the loss on the fine resolution $\mathcal{G}_{1}=\mathcal{G}$, only a few fine graph approximations are needed, and most of the computations are done on coarse, small graphs. Our method is summarized in Algorithm 2, and naturally suits inductive learning tasks, where multiple graphs are available.
@@
4.4 A Simple Model Problem
@@
Refer to caption Refer to caption Refer to caption
To explore this further, we train a four-layer GCN on a dataset of 100 rods in varying orientations and evaluate different coarsening methods for constructing a reduced graph. Table 1 presents the loss values for various coarsening strategies along with the corresponding edge counts. The results highlight a crucial aspect of our approach: methods that fail to increase connectivity prior to pooling (i.e., with $p=1$) lead to significant variance in loss across different coarsened graphs. For this dataset, introducing additional connectivity before pooling successfully mitigates this issue while still reducing the graph size in terms of both nodes and edges.

### v3

@@ Abstract
Graph Neural Networks (GNNs) have become powerful tools for learning from graph-structured data, finding applications across diverse domains. However, as graph sizes and connectivity increase, standard GNN training methods face significant computational and memory challenges, limiting their scalability and efficiency. In this paper, we present a novel framework for efficient multiscale training of GNNs. Our approach leverages hierarchical graph representations and subgraphs, enabling the integration of information across multiple scales and resolutions. By utilizing coarser graph abstractions and subgraphs, each with fewer nodes and edges, we significantly reduce computational overhead during training. Building on this framework, we propose a suite of scalable training strategies, including coarse-to-fine learning, subgraph-to-full-graph transfer, and multiscale gradient computation. We also provide some theoretical analysis of our methods and demonstrate their effectiveness across various datasets and learning tasks. Our results show that multiscale training can substantially accelerate GNN training for large scale problems while maintaining, or even improving, predictive performance.
@@ 1 Introduction
Large-scale graph networks arise in diverse fields such as social network analysis Kipf & Welling (2016); Defferrard et al. (2016), recommendation systems Tang et al. (2018), and bioinformatics Jumper et al. (2021), where relationships between entities are crucial to understanding system behavior Abadal et al. (2021). These problems involve processing large graph datasets, necessitating scalable solutions. Some examples of those problems include identifying community structures in social networks Hamilton et al. (2017), optimizing traffic flow in smart cities Geng et al. (2019), or studying protein-protein interactions in biology Xu et al. (2024). In all these problems, efficient processing of graph data is required for practical reasons.
In recent years, many Graph Neural Networks (GNNs) algorithms have been proposed to solve large graph problems, such as GraphSage Hamilton et al. (2017) and FastGCN Chen et al. (2018). However, despite their success, when the involved graph is of a large scale, with many nodes and edges, it takes a long time and many resources to train the network Galmés et al. (2021).
@@
(a) Random: Fine
(b) Random: Coarse
(c) Topk: Fine
(d) Topk: Coarse
Figure 1: Graph coarsening examples. Left two: random pooling; Right two: Topk pooling (highest-degree nodes selected). Orange nodes indicate selected coarse nodes.
@@
Figure 2: Coarse graph using subgraph pooling. Orange nodes indicate selected coarse nodes; the red node is the root for the ego-network.
@@
We demonstrate the generation of a coarse graph using random pooling and Topk pooling (which is based on the nodes with high degrees in the graph) in Figure 1. As for choosing a subgraph, this can be implemented using node-based selection methods, such as ego-networks (Gupta et al., 2014) or node marking and deletion (Frasca et al., 2022), which offer straightforward ways to extract subgraphs based on specific nodes of interest. We show an example of subgraph coarsening in Figure 2.
@@
While coarsening and subgraph methods have been in use for various purposes, from improved architectures Gao & Ji (2019), to enhanced expressiveness Bevilacqua et al. (2022; 2024), to the best of our knowledge, they were not considered in the context of efficient training of GNNs. Therefore, in this work, we explore three novel training mechanisms to design surrogate processes that approximate the original training problem with a fraction of the computational cost, using a multiscale approach. All the methods attempt to use the full graph with $n$ nodes as little as possible, and focus on training the network using coarse-scale graphs, thus reducing computational cost. The three methods, detailed in Section 4, are as follows:
  • Coarse-to-Fine. A coarse-to-fine algorithm approximates the optimization problem on smaller, pooled graphs. If the coarse-scale solution is similar to the fine-scale one, it can serve as an effective initialization for the fine-scale optimization process. Since gradient-based methods converge faster when initialized near the solution (Arora et al., 2018; Boyd & Vandenberghe, 2004), this approach reduces the overall number of iterations required.
  • Sub-to-Full. This approach leverages a sequence of subgraphs that progressively grow to the original fine-scale graph. Like coarsening, each subgraph is significantly smaller than the original graph, enabling faster optimization process. Subgraphs can be generated based on node distances (if available) or ego-network methods Gupta et al. (2014).
@@
The main contribution of this paper is the design of novel graph multiscale optimization algorithms to obtain computationally efficient GNN training methods that retain the performance of standard gradient-based training of GNNs. To demonstrate their effectiveness, we benchmark the obtained performance and reduce computational costs on multiple graph learning tasks, from transductive node classification to inductive tasks like shape segmentation.
@@ 2 Related Work
Graph pooling can be performed in various ways, as reviewed in Liu et al. (2022). Popular pooling techniques include DiffPool (Ying et al., 2018) MaxCutPool (Abate & Bianchi, 2024), graph wavelet compression Eliasof et al. (2023), as well as Independent-Set Pooling (Stanovic et al., 2024), all aiming to produce a coarsened graph that bears similar properties to the original graph. Wang et al. (2020) and Wang et al. (2024) introduce the concept of subgraph pooling to GNNs. Bianchi & Lachi (2023) study the expressive power of pooling in GNNs. Here, we use graph pooling mechanisms, as illustrated in Figure 1Figure 2, as core components for efficient training.
@@ 3 Preliminaries and Notations
We now provide related graph-learning notations and preliminaries that we will use throughout this paper. Let us consider a graph defined by ${\cal G=(\cal V,\cal E)}$ where $\cal V$ is a set of $n$ nodes and $\cal E$ is a set of $m$ edges. Alternatively, a graph can be represented by its adjacency matrix ${\bf A}$, where the $(i,j)$-th entry is 1 if an edge between nodes $i$ and $j$ exists, and 0 otherwise. We assume that each node $i$ is associated with a feature vector with $c$ channels, ${\bf x}_{i}\in\mathbb{R}^{c}$, and ${\bf X}=[{\bf x}_{0}\ldots,{\bf x}_{n-1}]\in\mathbb{R}^{n\times c}$ is the node features matrix. We denote by ${\bf y}_{i}\in\mathbb{R}^{d}$ the $i$-th node label, and by ${\bf Y}=[{\bf y}_{0},\ldots,{\bf y}_{n-1}]\in\mathbb{R}^{n\times d}$ the label matrix.
@@
For instance, in the well-known Graph Convolution Network (GCN) (Kipf & Welling, 2016), each layer is defined using Equation (1), such that ${\bf X}^{(l+1)}$ is the output of the $l$-th layer and serves as the input for the $(l+1)$-th layer. For a network with $L$ layers, we denote by ${\boldsymbol{\theta}}=\{{\bf W}^{(1)},\ldots,{\bf W}^{(L)}\}$ the collection of all learnable weight matrices within the network.
@@ 4 Efficient Training Framework for GNNs
${\bf Y}_{r}\approx f({\bf X}_{r},{{\cal G}_{r}};{\boldsymbol{\theta}}),\quad r=1,\ldots,R$ (3)
where $f$ is a GNN and ${\cal G}_{r}$ is a graph with $n_{r}$ nodes, where $n=n_{1}>n_{2}>\ldots>n_{R}$. Each graph ${\cal G}_{r}$ is obtained by pooling the fine graph $\mathcal{G}$, using a smaller set of nodes compared to the previous (finer) graph $\mathcal{G}_{r-1}$. ${\bf X}_{r}$ is the node features matrix of the graph $\mathcal{G}_{r}$, and ${\bf Y}_{r}$ is the corresponding label matrix. In what follows, we explore different ways to obtain ${\cal G}_{r},{\bf X}_{r}$ and ${\bf Y}_{r}$ from the original resolution variables.
@@ 4.1 Coarse-to-Fine Training
While Algorithm 1 is relatively simple, it requires obtaining the $r$-th surrogate problem by coarsening the graph, as well as a stopping criterion for the $r$-th optimization problem. Generally, in our coarse-to-fine method, we use the following graph coarsening scheme:
@@
where ${\bf P}_{r}\in\mathbb{R}^{n\times n_{r}}$ is a binary injection matrix that maps the fine graph with $n$ nodes to a smaller graph with $n_{r}$ nodes. That is, we choose a submatrix of the $p$-th power of the adjacency matrix ${\bf A}$. We use ${\bf X}_{r}={\bf P}_{r}^{T}{\bf X}$ as the coarse graph node features matrix and ${\bf Y}_{r}={\bf P}_{r}^{T}{\bf Y}$ as the label matrix of the coarse graph.
Choosing the non-zero entries of the matrix ${\bf P}_{r}$, which are the nodes that remain in the graph after pooling, can be implemented in various ways. For example, ${\bf P}_{r}$ can be obtained by sampling random nodes of the fine graph $\mathcal{G}$, as illustrated in Figure 1 (left), or with some desired pooling logic, such as choosing nodes of the highest degree in the graph, which is a method often described as $Topk$ pooling (Huang et al., 2015) and we present an illustration of it in Figure 1 (right). The power $p$ in Equation (5) controls the connectivity of the graph, where $p=1$ is the original graph adjacency matrix, and using a value $p>1$ will enhance the connectivity of the graph prior to the coarsening step.
In Algorithm 1, we apply GNN layers such as in Equation (1), using the coarsened feature matrix ${\bf X}_{r}$ and its corresponding adjacency matrix ${\bf A}_{r}$, that is based on a desired pooling ratio. We note that there is no re-initialization of the weight matrix ${\bf W}$ when transferring from coarse grids to finer ones, allowing a continuous learning process along the steps of the algorithm. As we show later in our experiments in Section 6, our coarse-to-fine training approach, summarized in Algorithm 1, is not restricted to a specific GNN architecture or a pooling technique.
@@ 4.2 Sub-to-Full Training
A subgraph can be defined in a few possible manners. One is by choosing a root node $j$, and generating a subgraph centered around that node. If the graph $\cal{G}$ has some geometric features, meaning each input node corresponds to a point in the Euclidean space $\mathbb{R}^{d}$, one might consider constructing a subgraph with $n_{r}$ nodes to be the $n_{r}$ nodes that are closest to the chosen center node in terms of Euclidean distance norm. Otherwise, we can consider a subgraph created using an ego-network Gupta et al. (2014) of $k-hop$ steps, such that we consider all neighbors that are $k$ steps away from the central node. An illustration of this method is presented in Figure 2. The number of hops, $k$, is a hyperparameter of the algorithm process, and we increase it as we get from coarse levels to fine ones within the series of graph $\{\mathcal{G}_{r}\}_{r=1}^{R}$.
@@ 4.3 Multiscale Gradients Computation
${\mathbb{E}}\,\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})={\mathbb{E}}\,\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})+{\mathbb{E}}\,\left(\mathcal{L}^{(R-1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})\right)+...+{\mathbb{E}}\,\left(\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(2)}_{s}({\boldsymbol{\theta}})\right),$ (8)
@@
${\mathbb{E}}\,\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})$$\approx$ $\textstyle{{\frac{1}{M_{R}}}\sum\,\mathcal{L}^{(R)}_{j}({\boldsymbol{\theta}})}+\textstyle{{\frac{1}{M_{R-1}}}\sum\,\left(\mathcal{L}^{(R-1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(R)}_{s}({\boldsymbol{\theta}})\right)}+...$
$+\textstyle{{\frac{1}{M_{1}}}\sum\,\left(\mathcal{L}^{(1)}_{s}({\boldsymbol{\theta}})-\mathcal{L}^{(2)}_{s}({\boldsymbol{\theta}})\right)},$
where $M_{R}>M_{R-1}>M_{1}$. This implies that for the computation of the loss on the fine resolution $\mathcal{G}_{1}=\mathcal{G}$, only a few fine graph approximations are needed, and most of the computations are done on coarse, small graphs. Our method is summarized in Algorithm 2, and we provide an illustration of the computation process in Figure 3. Naturally, this method suits inductive learning tasks, where multiple graphs are available.
@@
Figure 3: Illustration of the Multiscale Gradient Computation algorithm introduced in Section 4.3.
4.4 A Motivating Example
@@
Refer to caption Refer to caption Refer to caption
To explore this further, we train a four-layer GCN Kipf & Welling (2016) on a dataset of 100 rods in varying orientations and evaluate different coarsening methods for constructing a reduced graph. Table 1 presents the loss values for various coarsening strategies along with the corresponding edge counts. The results highlight another aspect of our approach: methods that do not increase connectivity prior to pooling (i.e., with $p=1$) may, in certain situations, lead to significant variance in loss across different coarsened graphs. For this dataset, introducing additional connectivity before pooling successfully mitigates this issue while still reducing the graph size in terms of both nodes and edges.
To further evaluate the capabilities of our multiscale methods on this synthetic dataset, Appendix B presents the full multiscale training results and compares them with standard fine-grid training. Additionally, we construct a second synthetic dataset derived from the MNIST image dataset LeCun et al. (1998). Results on both datasets demonstrate that our methods effectively capture global graph features while enabling efficient training.
4.5 Theoretical Properties of Reduced Graphs
At the base of all the above discussed techniques, the core is the ability of the coarse graph to represent the fine graph. To this end, consider the model problem
${\ell}({\boldsymbol{\Theta}})={\frac{1}{2N}}\|\sigma({\bf A}{\bf X}{\boldsymbol{\Theta}})-{\bf Y}\|^{2},$ (11)
and assume that $\sigma$ is the identity mapping. One way to analyze the solution of the coarse-scale problem is through sketching (see Woodruff (2014)). In sketching, the original problem is replaced by
${\ell}_{c}({\boldsymbol{\Theta}})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf X}{\boldsymbol{\Theta}}-{\bf P}^{\top}{\bf Y}\|^{2}$ (12)
where ${\bf P}^{\top}$ is a sketching matrix. We establish the following result:
Theorem 4.1.
Let ${\boldsymbol{\Theta}}^{*}$ be the solution of equation (11) and let ${\boldsymbol{\Theta}}_{C}^{*}$ be the solution of equation (12). Let ${\bf P}^{\top}$ be a sparse subspace embedding with $O(\frac{c^{2}}{\epsilon})$ rows, where $c$ is the size of the feature vectors (number of graph channels), that injects any vector into a subset of indices, which we denote by $C$. According to this subset, denote ${\bf A}=\begin{bmatrix}{\bf A}_{CC}&{\bf A}_{CF}\\ {\bf A}_{FC}&{\bf A}_{FF}\end{bmatrix}$, and ${\bf X}=\begin{bmatrix}{\bf X}_{C}\\ {\bf X}_{F}\end{bmatrix}$. Then if $\sigma=Id$, there exists $\|{\bf R}\|<\|{\bf A}_{CF}{\bf X}_{F}{\boldsymbol{\Theta}}_{C}^{*}\|$ such that with probability .99:
${\ell}_{c}({\boldsymbol{\Theta}}_{C}^{*})={\frac{1}{2N}}\|{\bf P}^{\top}{\bf A}{\bf P}{\bf P}^{\top}{\bf X}{\boldsymbol{\Theta}}_{C}^{*}-{\bf P}^{\top}{\bf Y}\|^{2}\leq{\frac{1\pm\epsilon}{2N}}\left(\|{\bf A}{\bf X}{\boldsymbol{\Theta}}^{*}-{\bf Y}\|^{2}+\|{\bf R}\|^{2}\right)$ (13)
The proof is provided in Appendix A.